20. Still using the same Demand and Supply curves that can be represented by:     Qd =...

20.

Still using the same Demand and Supply curves that can be represented by:     Qd = 90-2P; and Qs = 3P.   Suppose price is at $15, there will be ……..    as many as ……..units.

shortage; 60 units
surplus; 45 units
shortage; 15 units
surplus; 15 units
Answer #1

Qd = 90-2P

Qs= 3P

At P=$15,

Qd = 90-2(15)= 90-30 = 60 units

Qs = 3(15)= 45 units

Because Qd>Qs , this implies that at a price of $15 , there will be a shortage as many as (60-45)= 15 units.

Hence, option(C) is correct.